# Deep Learning for Clouds and Cloud Shadow Segmentation in Methane Satellite and Airborne Imaging Spectroscopy

### v2

@@ III Data
This study utilizes calibrated and georeferenced Level 1B (L1B) hyperspectral data from MethaneAIR and MethaneSAT Conway_2024 , with cloud labels derived from Level2 (L2) derived quantities [^1] To download the data and code visit https://doi.org/10.7910/DVN/IKLZOJ ccm2024methaneair . The L2 products consist of the retrieved fitted parameters obtained from the best fit of a radiative transfer model to the measured L1B spectra. The main products from the CH4 spectrometer are the retrieved CO2 and CH4 vertical column densities (VCDs), and the main product from the O2 spectrometer is the retrieved surface pressure.
@@ III-A MethaneAIR
MethaneAIR ccm2024methaneair ; chulakadabba2023methane ; Guanter_2024 ; Warren_2024 is an airborne simulator for the MethaneSAT satellite mission, designed to capture high-resolution spectral imagery for atmospheric composition analysis. MethaneAIR flights are conducted at typically 40,000 ft above ground. MethaneAIR measurements have been conducted on both NCAR GV and Learjet-35 platforms, and over 100 flight hours have been completed across the US, observing major oil and gas production areas. Each individual MethaneAIR campaign maps roughly 100 km x 100 km regions in over 2 hours, which are resolved at high precision and high spatial resolution.
@@ IV Methodology
Our approach for cloud and shadow detection in hyperspectral satellite imagery comprises four main components: data preprocessing, semantic segmentation model development, a training procedure, and a results evaluation procedure. For MethaneSAT, we developed a post-processing component that allows us to evaluate our models despite its variable spatial dimensions. We first establish a preprocessing pipeline that handles missing values and implements strategic normalization steps to ensure consistent input features. Then, we perform a comparison of various semantic segmentation approaches spanning from traditional machine learning to advanced deep learning architectures. Finally, we detail a training procedure that addresses class imbalance and incorporates data augmentation techniques to enhance model generalization. For MethaneSAT, we present an alternative post-processing strategy that uses fixed-size overlapping patches to accommodate the varying input spatial dimensions.
IV-A Preprocessing Steps
To ensure high-quality input data and optimize model performance, we designed a rigorous preprocessing pipeline. The first step addresses missing values in the hyperspectral data. Any missing values (NaNs) within the dataset are imputed using the mean value along the spectral dimension for each affected pixel, ensuring continuity in the spectral signatures. Missing values may arise, for example, over very dark or very bright features on the landscape.
@@ IV-A Preprocessing Steps
Before passing the input data to any machine learning model, a batch-wise normalization is performed across all dimensions. In this step, each sample in the batch is normalized by subtracting the mean value calculated across all spectral bands and spatial dimensions (height and width) and dividing by the standard deviation computed across these same dimensions. This approach treats each hyperspectral data cube as a unified entity, normalizing it to have zero mean and unit variance across all its elements. This batch-sample normalization helps reduce the influence of overall intensity variations between different images while maintaining the relative relationships between spectral bands and spatial structures that are crucial to distinguish between clouds, shadows, and clear sky.
IV-B Semantic Segmentation Algorithms
To address the task of cloud and shadow identification in MethaneAIR hyperspectral imagery, we evaluated diverse machine learning methods: Iterative Logistic Regression (LR; Edisanter_2016 ; park_2023 ; Gautam_2023 ), Multilayer Perceptron (MLP; BENEDIKTSSON_1993 ; Yang_1999 ; Tian_1999 ; Taravat_2015 ; gomezchova_2017 ), U-Net Ronneberger2015 ; Jiao_2020 ; Wieland_2019 ; Miao_2022 ; wright_2024 ; Li_2022 ; Tan_2023 , and the Spectral Attention Network (SCAN), along with the combined models.
IV-B1 Iterative Logistic Regression (ILR)
We implement the Iterative Logistic Regression (ILR) approach introduced by park_2023 as a baseline model for our hyperspectral shadow detection system. This established method learns a compact spectral basis that effectively distinguishes between shadowed and non-shadowed spatial soundings through an iterative dimensionality reduction process.
Given an input tensor $X\in\mathbb{R}^{H\times W\times C}$, where $H$ and $W$ are spatial dimensions, and $C$ is the number of spectral channels, the algorithm takes each sounding spectrum $x_{i}\in\mathbb{R}^{C}$ in the dataset, compute its mean radiance $m_{i}$ and apply a normalization transformation:
$s_{i}=\log(x_{i}/m_{i})$
where $\log$ denotes the natural logarithm. This preprocessing step isolates the spectral shape features from overall brightness variations. The core algorithm iteratively extracts spectral components most relevant for cloud and shadow detection:
1. Using log spectra $s_{i}$ and their associated labels $y_{i}\in\mathcal{R}^{K}$, where K is the number of classes, the algorithm trains a logistic regression classifier to find spectral weights $w_{t}$ that separate each of the K classes.
@@ IV-B1 Iterative Logistic Regression (ILR)
3. This process repeats $z$ times until the classification performance (measured by F1-score) falls below a predefined threshold.
@@
Through this iterative process, we compute a low-dimensional representation with $z=23$ components (compared to the original 1024 and 1080 spectral bands for MethaneAIR and MethaneSAT respectively), which captures the most salient spectral features for shadow detection.
The learned representation $\beta_{i}\in\mathcal{R}^{z}$ is multiplied by a weight matrix $W_{class}\in\mathbb{R}^{z\times K}$ and the Softmax activation function is applied. This way, the probability of belonging to each class can be estimated as:
@@
$P(y_{i}|x_{i})=\text{Softmax}(W_{class}^{T}\beta_{i})$
IV-B2 Multilayer Perceptron (MLP)
@@
Building upon the spectral feature extraction approach of ILR, we implement a Multilayer Perceptron (MLP; werbos_1970 ) model that processes hyperspectral data on a pixel-wise basis. Given an input tensor $X\in\mathbb{R}^{H\times W\times C}$, where $H$ and $W$ are spatial dimensions, and $C$ is the number of spectral channels (1024 for MethaneAIR and 1080 for MethaneSAT), the MLP treats each pixel spectrum $x_{i}\in\mathbb{R}^{C}$ independently.
@@
Our MLP implementation processes each pixel’s complete spectral signature independently through a neural network architecture optimized for the dimensionality of our data. The network architecture consists of an input layer with $C$ nodes, followed by two hidden layers with 20 nodes each, and an output layer with $K$ nodes corresponding to our target classes. Mathematically, for each pixel spectrum $x_{i}$, the probability of belonging to each class can be computed as:
$h^{(1)}_{i}=\phi(W^{(1)}x_{i}+b^{(1)})$
$h^{(2)}_{i}=\phi(W^{(2)}h^{(1)}_{i}+b^{(2)})$
$P(y_{i}|x_{i})=\text{Softmax}(W^{(3)}h^{(2)}_{i}+b^{(3)})$
@@
where $W^{(1)}\in\mathbb{R}^{20\times C}$, $W^{(2)}\in\mathbb{R}^{20\times 20}$, and $W^{(3)}\in\mathbb{R}^{K\times 20}$ are the weight matrices, $b^{(1)}\in\mathbb{R}^{20}$, $b^{(2)}\in\mathbb{R}^{20}$, and $b^{(3)}\in\mathbb{R}^{K}$ are the bias vectors, and $\phi(\cdot)$ represents the Rectified Linear Unit (ReLU) activation function.
@@
Unlike the ILR approach, which explicitly reduces dimensionality through iterative feature extraction, the MLP learns a direct non-linear mapping from the full spectral signature to class probabilities. This allows the network to capture complex spectral patterns without the need for explicit feature engineering. The compact architecture (only 20 nodes per hidden layer) maintains computational efficiency while providing sufficient capacity to model non-linear relationships between spectral bands that are relevant for cloud and shadow detection.
IV-B3 U-Net Architecture
We adopt the U-Net architecture (Ronneberger2015, ) for hyperspectral semantic segmentation due to its demonstrated effectiveness in dense prediction tasks through its symmetric encoder-decoder structure with skip connections. U-Net has been shown to perform well in hyperspectral segmentation tasks by efficiently capturing both spatial and spectral features across different scales (Wei_2023, ; ruzicka_2023, ).
Given an input hyperspectral image $X\in\mathbb{R}^{B\times C\times H\times W}$, where $B$ is the batch size, $C$ is the number of spectral channels, and $H$ and $W$ are spatial dimensions, our U-Net processes the data through the following components:
@@ IV-B3 U-Net Architecture
Output Layer: The final layer produces a tensor with dimensions $B\times K\times H\times W$, where $K$ is the number of classes. A softmax function is applied to generate a probability distribution over classes for each pixel:
$P(y|X)=\text{softmax}(f_{U-Net}(X))$
@@
IV-B4 Spectral Channel Attention Network
@@
Attention mechanisms have been widely adopted in computer vision, their application to spectral band selection in hyperspectral cloud/shadow segmentation presents special challenges and opportunities. Unlike natural images where spatial attention focuses on semantic objects, spectral attention specifically addresses the band selection problem inherent in hyperspectral analysis. The key insight is that different spectral regions provide varying levels of discrimination between clouds/shadows and surface materials, particularly at boundary regions where spectral mixing occurs.
@@
We reformulate the classical attention mechanism from bahdanau_2014 as a Spectral Channel Attention Network (SCAN), adapting channel-wise attention for hyperspectral band selection in cloud and shadow detection tasks. This approach directly addresses the boundary delineation problem by dynamically weighting the importance of different spectral bands based on their discriminative power for cloud/shadow detection. Rather than treating all spectral bands equally as in conventional approaches, our method learns to emphasize wavelengths that are most informative for distinguishing clouds/shadows from underlying surfaces, particularly in spectrally ambiguous transition zones. The model consists of two main components: a spectral attention module and a pixel-wise classification framework.
Spectral Attention Module: Given an input hyperspectral image $X\in\mathbb{R}^{B\times H\times W\times C}$, where $B$ is the batch size, $H$ and $W$ are spatial dimensions, and $C$ is the number of spectral channels, we apply a channel-wise attention mechanism that learns band-specific importance weights. The attention weights $\alpha$ are computed as:
$\alpha=\sigma(W_{2}\text{ReLU}(W_{1}\bar{x}))$
where $\bar{x}\in\mathbb{R}^{C}$ is the spatially averaged input, $W_{1}\in\mathbb{R}^{C/16\times C}$ and $W_{2}\in\mathbb{R}^{C\times C/16}$ are learnable parameters, and $\sigma$ is the sigmoid activation function. The attended features are obtained through:
$X_{att}=X\odot\alpha$
where $\odot$ represents channel-wise multiplication broadcasting the attention weights across all spatial locations. Intuitively, $X_{att}\in\mathbb{R}^{B\times H\times W\times C}$ has the same input dimension and represents a spectrally-weighted version of the input that emphasizes the most discriminative bands for cloud/shadow detection while suppressing less informative spectral regions.
@@
Classification Framework: The attended features are passed through a fully-connected neural network as follows:
$f_{MLP}(X)=W_{n}(\phi(W_{n-1}(...\phi(W_{1}X_{att})))),$
@@
and the final probabilities are obtained through a softmax activation function:
$P(y|X)=\text{Softmax}(f_{MLP}(X_{att})).$
IV-B5 Combined Models
To leverage the complementary strengths of spatial and spectral approaches, we developed two ensemble methods that combine the U-Net and Spectral Channel Attention Network (SCAN) models: Combined MLP and Combined CNN.

### v3

@@ III Data
This study utilizes calibrated and georeferenced Level 1B (L1B) hyperspectral data from MethaneAIR and MethaneSAT Conway et al. (2024), with cloud labels derived from Level2 (L2) derived quantities [^1] To download the data and code visit https://doi.org/10.7910/DVN/IKLZOJ Chan Miller et al. (2024). The L2 products consist of the retrieved fitted parameters obtained from the best fit of a radiative transfer model to the measured L1B spectra. The main products from the CH4 spectrometer are the retrieved CO2 and CH4 vertical column densities (VCDs), and the main product from the O2 spectrometer is the retrieved surface pressure.
@@ III-A MethaneAIR
MethaneAIR Chan Miller et al. (2024); Chulakadabba et al. (2023); Guanter et al. (2025); Warren et al. (2024) is an airborne simulator for the MethaneSAT satellite mission, designed to capture high-resolution spectral imagery for atmospheric composition analysis. MethaneAIR flights are conducted at typically 40,000 ft above ground. MethaneAIR measurements have been conducted on both NCAR GV and Learjet-35 platforms, and over 100 flight hours have been completed across the US, observing major oil and gas production areas. Each individual MethaneAIR campaign maps roughly 100 km x 100 km regions in over 2 hours, which are resolved at high precision and high spatial resolution.
@@
While the L2-derived cloud flags provide a physically consistent and operationally relevant basis for training labels, we recognize the potential for propagating systematic biases from the screening algorithm. To mitigate this, we implemented a rigorous manual curation process. All 508 scenes were visually inspected to verify label quality and consistency with the observed radiance patterns. We excluded scenes exhibiting: (1) ambiguous cloud/shadow boundaries where L2 flags showed inconsistencies with visual inspection, (2) systematic artifacts such as dark surface misclassification, and (3) full cloud coverage or completely clear conditions that provide limited training diversity. This curation ensures that our training data represents high-confidence examples where L2-derived labels align with true atmospheric conditions.
@@ III-B MethaneSAT
Similar to MethaneAIR, we manually curated the MethaneSAT dataset through visual inspection of all available scenes. Specifically, we excluded: (1) scenes containing water bodies, (2) images with full cloud coverage that provide limited training value, and (3) any cases where visual inspection revealed systematic inconsistencies between the L2-derived labels and the apparent cloud/shadow patterns in the radiance imagery. This quality control process retained 262 high-quality scenes spanning diverse geographic regions, surface types, and atmospheric conditions, ensuring our training data represents reliably labeled examples rather than systematic artifacts of the L2 screening algorithm.
@@ IV Methodology
Our approach for cloud and shadow detection in hyperspectral satellite imagery comprises three main components: preprocessing, semantic segmentation, and a training procedure. Additionally, an evaluation framework is developed to compare model performances. For MethaneSAT, we also developed a post-processing component that handles variable spatial dimensions through patch-based inference. An overview of our approach is displayed in Figure 4
We first establish a formal problem statement, then systematically evaluate diverse architectures spanning spectral-only methods (ILR, MLP), spatial methods (U-Net), and spectral channel attention mechanism (SCAN). This comprehensive comparison reveals the complementary strengths of different approaches and motivates our best-performing ensemble architectures (Combined MLP/CNN).
IV-A Problem Formulation
Given a hyperspectral data cube $\mathbf{X}\in\mathbb{R}^{H\times W\times C}$, where $H$ and $W$ represent spatial dimensions and $C$ denotes spectral channels ($C=1024$ for MethaneAIR, $C=1080$ for MethaneSAT), our objective is to learn a mapping function $f:\mathbb{R}^{H\times W\times C}\rightarrow\mathbb{R}^{H\times W\times K}$ that assigns each spatial sounding to one of $K$ classes (background, cloud, shadow, and optionally dark surface for MethaneAIR, where $K=4$; background, cloud, shadow for MethaneSAT, where $K=3$).
Formally, we seek to minimize the expected loss:
$\mathcal{L}(\mathbf{W})=\mathbb{E}_{(\mathbf{X},\mathbf{Y})\sim\mathcal{D}}\left[\ell(f_{\mathbf{W}}(\mathbf{X}),\mathbf{Y})\right]$ (1)
where $\mathbf{W}$ represents learnable parameters, $\mathbf{Y}\in\{0,1\}^{H\times W\times K}$ is the one-hot encoded ground truth segmentation mask, $\ell$ is the weighted cross-entropy loss (detailed in Section V), and $\mathcal{D}$ is the distribution of hyperspectral scenes.
Figure 4: Methodology overview for cloud and shadow detection in hyperspectral methane monitoring data. The pipeline processes input hyperspectral cubes X $\in\mathbb{R}^{H\times W\times C}$ through preprocessing steps (missing value imputation, spectral normalization) before semantic segmentation. We evaluate six model architectures: spectral-only methods (ILR, MLP), spatial convolution (U-Net), spectral attention (SCAN), and ensemble fusion approaches (Combined MLP/CNN) that integrate predictions from frozen U-Net and SCAN models (dashed arrows). All models are optimized using weighted cross-entropy loss $\ell(f_{\mathcal{W}}(\mathbf{X}),\mathbf{Y})$ to produce pixel-wise cloud and shadow masks Y $\in\{0,1\}^{H\times W\times K}$ with $K=4$ classes (MethaneAIR) or $K=3$ classes (MethaneSAT).
IV-B Preprocessing Steps
To ensure high-quality input data and optimize model performance, we designed a rigorous preprocessing pipeline. The first step addresses missing values in the hyperspectral data. Any missing values (NaNs) within the dataset are imputed using the mean value along the spectral dimension for each affected spatial sounding, ensuring continuity in the spectral signatures. Missing values may arise, for example, over very dark or very bright features on the landscape.
@@ IV-B Preprocessing Steps
Before passing the input data to any machine learning model, a batch-wise normalization is performed across all dimensions. In this step, each sample in the batch is normalized by subtracting the mean value calculated across all spectral bands and spatial dimensions (height and width) and dividing by the standard deviation computed across these same dimensions. This batch-sample normalization helps reduce the influence of overall intensity variations between different images while maintaining the relative relationships between spectral bands and spatial structures that are crucial to distinguish between clouds, shadows, and clear sky.
IV-C Semantic Segmentation Algorithms
We evaluate six distinct architectures for cloud and shadow segmentation, ranging from conventional spectral-based methods to advanced deep learning approaches: Iterative Logistic Regression (LR; Lo and Ientilucci (2016); Park et al. (2023); Gautam and Nadda (2023)), Multilayer Perceptron (MLP; J. A. BENEDIKTSSON and ERSOY (1993); Yang (1999); Tian et al. (1999); Taravat et al. (2015); Gómez-Chova et al. (2017)), U-Net Ronneberger et al. (2015); Jiao et al. (2020); Wieland et al. (2019); Miao et al. (2022); Wright et al. (2024); Li et al. (2022); Tan et al. (2023), and the Spectral Attention Network (SCAN), along with the combined models. This systematic comparison reveals the progression of capabilities and the complementary strengths of spatial versus spectral modeling.
IV-C1 Iterative Logistic Regression (ILR)
We implement the Iterative Logistic Regression (ILR) approach introduced by Park et al. (2023) as a baseline model for our hyperspectral shadow detection system. This established method learns a compact spectral basis that effectively distinguishes between shadowed and non-shadowed spatial soundings through an iterative dimensionality reduction process.
Given an input spectrum $\mathbf{x}_{i}\in\mathbb{R}^{C}$ for spatial sounding $i$, we first compute its mean radiance $m_{i}=\frac{1}{C}\sum_{c=1}^{C}x_{i,c}$ and apply a logarithmic normalization transformation:
$\mathbf{s}_{i}=\log\left(\mathbf{x}_{i}/m_{i}\right)$ (2)
where $\log$ denotes element-wise natural logarithm. This preprocessing step isolates spectral shape features from overall brightness variations, making the representation more robust to illumination changes. The core algorithm iteratively extracts spectral components most relevant for cloud and shadow detection:
1. Using log spectra $s_{i}$ and their associated labels $y_{i}\in\mathcal{R}^{K}$, where K is the number of classes, the algorithm trains a logistic regression classifier to find spectral weights $\mathbf{W}$ that separate each of the K classes.
@@ IV-C1 Iterative Logistic Regression (ILR)
3. This process repeats $z$ times until the classification performance (measured by F1-score) converges or falls below a predefined threshold.
@@
Through this iterative process, we compute a low-dimensional representation $\boldsymbol{\beta}_{i}\in\mathbb{R}^{z}$, with $z=23$ components (compared to the original 1024 and 1080 spectral bands for MethaneAIR and MethaneSAT respectively), which captures the most salient spectral features for shadow detection.
The learned low-dimensional representations $\boldsymbol{\beta}_{i}$ are used for to obtain output probabilities for each spatial sounding:
@@
$P(\mathbf{y}_{i}|\mathbf{x}_{i})=\text{softmax}\left(\mathbf{W}_{\text{class}}^{T}\boldsymbol{\beta}_{i}\right)$ (3)
where $\mathbf{W}_{\text{class}}\in\mathbb{R}^{z\times K}$ is trained via multinomial logistic regression on the compressed features.
IV-C2 Multilayer Perceptron (MLP)
@@
Building upon the spectral feature extraction approach of ILR, we implement a Multilayer Perceptron (MLP; Werbos (1970)) model that processes hyperspectral data on a pixel-wise basis. The MLP treats each spatial sounding $\mathbf{x}_{i}\in\mathbb{R}^{C}$ independently, learning a direct non-linear mapping from the full spectral signature to class probabilities without explicit dimensionality reduction.
@@
Our MLP implementation processes each spatial sounding’s complete spectral signature independently through a neural network architecture. The network architecture consists of an input layer with $C$ nodes, followed by two hidden layers with with $d_{1}=d_{2}=20$ nodes each, and an output layer with $K$ nodes corresponding to our target classes. Mathematically, for each spatial sounding $x_{i}$, the probability of belonging to each class can be computed as:
$\mathbf{h}^{(1)}_{i}$ $=\phi\left(\mathbf{W}^{(1)}\mathbf{x}_{i}+\mathbf{b}^{(1)}\right)$ (4)
$\mathbf{h}^{(2)}_{i}$ $=\phi\left(\mathbf{W}^{(2)}\mathbf{h}^{(1)}_{i}+\mathbf{b}^{(2)}\right)$ (5)
$P(\mathbf{y}_{i}|\mathbf{x}_{i})$ $=\text{softmax}\left(\mathbf{W}^{(3)}\mathbf{h}^{(2)}_{i}+\mathbf{b}^{(3)}\right)$ (6)
@@
where $\mathbf{W}^{(1)}\in\mathbb{R}^{20\times C}$, $\mathbf{W}^{(2)}\in\mathbb{R}^{20\times 20}$, and $\mathbf{W}^{(3)}\in\mathbb{R}^{K\times 20}$ are weight matrices; $\mathbf{b}^{(1)}\in\mathbb{R}^{20}$, $\mathbf{b}^{(2)}\in\mathbb{R}^{20}$, and $\mathbf{b}^{(3)}\in\mathbb{R}^{K}$ are bias vectors; and $\phi(\cdot)=\max(0,\cdot)$ is the Rectified Linear Unit (ReLU; Agarap (2018)) activation function.
@@
Unlike ILR, which explicitly reduces dimensionality through iterative feature extraction, the MLP learns a direct non-linear mapping from the full spectral signature to class probabilities. This allows the network to capture complex spectral patterns without the need for explicit feature engineering, potentially discovering discriminative spectral relationships not captured by linear projections.
IV-C3 U-Net Architecture
We adopt the U-Net architecture (Ronneberger et al., 2015) for its proven effectiveness in dense prediction tasks through its multi-scale spatial feature extraction. Unlike pixel-wise methods (ILR, MLP), U-Net captures hierarchical spatial features through its symmetric encoder-decoder structure with skip connections, enabling accurate boundary delineation essential for precise cloud and shadow detection (Wei et al., 2023; Růžička et al., 2023).
@@ IV-C3 U-Net Architecture
Output Layer: The final layer produces a tensor with dimensions $K\times H\times W$, where $K$ is the number of classes. A softmax function is applied to generate a probability distribution over classes for each spatial sounding:
$P(\mathbf{Y}|\mathbf{X})=\text{softmax}(f_{U-Net}(\mathbf{X}))$
@@
IV-C4 Spectral Channel Attention Network
@@
Attention mechanisms have been widely adopted in computer vision, but their application to spectral band selection in hyperspectral cloud/shadow segmentation presents special challenges and opportunities. Unlike natural images where spatial attention focuses on semantic objects, spectral attention specifically addresses the band selection problem inherent in hyperspectral analysis. The key insight is that different spectral regions provide varying levels of discrimination between clouds/shadows and surface materials, particularly at boundary regions where spectral mixing occurs.
@@
We reformulate the classical attention mechanism from Bahdanau et al. (2014) as a Spectral Channel Attention Network (SCAN), adapting channel-wise attention for hyperspectral band selection in cloud and shadow detection tasks. An overview of this model is displayed in Appendix -B Figure 13. This approach directly addresses the boundary delineation problem by dynamically weighting the importance of different spectral bands based on their discriminative power for cloud/shadow detection. Rather than treating all spectral bands equally as in conventional approaches, our method learns to emphasize wavelengths that are most informative for distinguishing clouds/shadows from underlying surfaces, particularly in spectrally ambiguous transition zones. The model consists of two main components: a spectral attention module and a pixel-wise classification framework.
Spectral Attention Module. Given input $\mathbf{X}\in\mathbb{R}^{H\times W\times C}$, we first compute spatially-pooled features that summarize global spectral characteristics:
$\bar{\mathbf{x}}=\text{GlobalAveragePooling}(\mathbf{X})\in\mathbb{R}^{C}$ (7)
This spatial aggregation is motivated by the observation that spectral discriminability is a global property (e.g., the 1640 nm band is informative for cloud detection everywhere in the scene), not a local, position-dependent characteristic. The pooled features are then passed through a multilayer perceptron (MLP) to generate attention weights:
$\mathbf{z}$ $=\text{ReLU}\left(\mathbf{W}_{1}\bar{\mathbf{x}}+\mathbf{b}_{1}\right)\quad\text{where }\mathbf{W}_{1}\in\mathbb{R}^{\frac{C}{r}\times C}$ (8)
$\boldsymbol{\alpha}$ $=\sigma\left(\mathbf{W}_{2}\mathbf{z}+\mathbf{b}_{2}\right)\quad\text{where }\mathbf{W}_{2}\in\mathbb{R}^{C\times\frac{C}{r}}$ (9)
where $\sigma(\cdot)$ is the sigmoid function, $r=16$ is the reduction ratio controlling bottleneck dimensionality, and $\boldsymbol{\alpha}\in[0,1]^{B\times C}$ represents learned channel importance weights. Values near 1 amplify discriminative bands; values near 0 suppress uninformative bands.
The attended features are obtained via channel-wise multiplication ($\boldsymbol{\alpha}$ across spatial dimensions):
$\mathbf{X}_{\text{att}}=\mathbf{X}\odot\alpha$ (10)
@@
Classification Framework: The attended features are passed through a fully-connected neural network with $N$ layers as follows:
$f_{MLP}(\mathbf{X})=\mathbf{W^{\prime}}_{N}(\phi(\mathbf{W^{\prime}}_{N-1}(...\phi(\mathbf{W^{\prime}}_{1}X_{att})))),$
@@
where $\mathbf{W^{\prime}}_{i}$ are the classifier learnable weight matrices, $\phi$ denotes ReLU activation function. The final probabilities are obtained through a softmax activation function:
$P(\mathbf{Y}|\mathbf{X})=\text{Softmax}(f_{MLP}(\mathbf{X}_{att})).$
Note that the spectral attention module parameters $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$, along with the classifier parameters $\mathbf{W^{\prime}}_{i}$, are jointly optimized during training.
IV-C5 Combined Models
To leverage the complementary strengths of spatial (U-Net) and spectral (SCAN) approaches, we develop ensemble architectures that fuse their predictions. We explore two fusion strategies: MLP-based fusion (pixel-wise combination) and CNN-based fusion (spatially-aware combination).